OpenStudy (anonymous):

–6p + 7 = 3(2p – 3) – 4(–10 + 4p)

OpenStudy (anonymous):

distribute the 3 and 4 to get rid of the parenthesis

OpenStudy (anonymous):

-6p+7=6p-9+40-8p -6p=6p+24-8p -6p=-2p+24 -4p=24 p=-6